OpenStudy (sleepyjess):

Proving trig equations (sin x)(tan x cos x - cot x cos x) = 1 - 2 cos^2 x

OpenStudy (unklerhaukus):

On the left hand side of the equation use: \[\tan x=\frac{\sin x}{\cos x}\] and: \[\cot x=\frac1{\tan x}=\frac{\cos x}{\sin x}\]

OpenStudy (sleepyjess):

\((sin x)(\dfrac{sinx}{cosx}cosx-\dfrac{cosx}{sinx}cosx)=1-2soc^2x\)

OpenStudy (nincompoop):

uh try to fix what is inside the second parenthesis, see how you can manipulate it

OpenStudy (unklerhaukus):

good, now you can simplify this part \[\frac{\sin x}{\cos x}{\cos x}\]

OpenStudy (sleepyjess):

Wouldn't \(\dfrac{sinx}{cosx}\) and \(\dfrac{cosx}{sinx}\) cancel?

OpenStudy (nincompoop):

if you're multiplying and it does not cancel, it gives you 1

OpenStudy (sleepyjess):

Oh ok

OpenStudy (sleepyjess):

\((sin x)(1-\dfrac{cosx}{sinx}cosx)=1-2cos^2x\)

OpenStudy (nincompoop):

whoa whoa whoa how is that possible?

OpenStudy (unklerhaukus):

\[(\sin x)\left(\frac{\sin x}{\cos x}{\cos x}-\frac{\cos x}{\sin x}\cos x\right)=1-2\cos^2x\\ (\sin x)\left(\frac{\sin x}{\cancel{\cos x}}{\cancel{\cos x}}-\frac{\cos x}{\sin x}\cos x\right)=1-2\cos^2x\\\vdots\]

OpenStudy (sleepyjess):

Oh so \((sin x)(sinx-\dfrac{cosx}{sinx}cosx)=1-2cos^2x\) right?

OpenStudy (nincompoop):

you're not done yet!

OpenStudy (unklerhaukus):

that's better, now distribute the (sin x) into the big brackets

OpenStudy (sleepyjess):

Ok, \((sinx^2-\dfrac{cosx}{sinx}cosx~sinx)=1-2cos^2x\). Then \(\dfrac{cosx}{sinx}cosx~sinx\) cancels to become cosx*cosx?

OpenStudy (unklerhaukus):

yeah and cos x * cos x = cos^2 x

OpenStudy (sleepyjess):

Ok, so now it is \(sin^2x-cos^2x=1-2cos^2x\)?

OpenStudy (unklerhaukus):

good

OpenStudy (unklerhaukus):

now what do you get if you add cos^2x to both sides?

OpenStudy (sleepyjess):

\(\sf sin^2x=1\)

OpenStudy (unklerhaukus):

not quite

OpenStudy (sleepyjess):

Oh, no, \(\sf sin^2x=1-cos^2x\)

OpenStudy (unklerhaukus):

can you rearrange this to the one trig identity you can remember?

OpenStudy (sleepyjess):

\(\sf sin^2x+cos^2x=1\)?

OpenStudy (unklerhaukus):

yep, !

OpenStudy (sleepyjess):

:D

Nnesha (nnesha):

wait but when we have to prove both sides we can solve only one side mean to say cannot move left to right or right to left anything right ??

OpenStudy (unklerhaukus):

Why is that a condition of a proof?

Nnesha (nnesha):

:D nvm my teacher said that

OpenStudy (anonymous):

Nnesha wht ur teacher said?

OpenStudy (mathmath333):

my teacher also used to say that

ganeshie8 (ganeshie8):

your both teachers are good they are guiding you how to write a neat looking proof :)

OpenStudy (unklerhaukus):

What is wrong with using equations?

OpenStudy (nincompoop):

nothing it is an equation from beginning to end LAUGHING OUT LOUD

ganeshie8 (ganeshie8):

nothing wrong adding stuff to both sides and messing with both sides looks messy.. thats all..

OpenStudy (mathmath333):

it doesnt change much when u use both sides to prove , the proof remains true

Nnesha (nnesha):

^^^ lol yeah thanks to all of you but btw if i do same thing she probably gonna take offff ppppoints :D:(

ganeshie8 (ganeshie8):

first 5 minutes of this video https://www.youtube.com/watch?v=NuGDkmwEObM

OpenStudy (mathmath333):

it says "well organized".

ganeshie8 (ganeshie8):

it says below strategy doesnt work always but i can't think of a simple example where it could fail at the moment : A = B ..=.. ..=.. 1 = 1

ganeshie8 (ganeshie8):

im sure this shouldn't matter for almost all the proofs in trig

Nnesha (nnesha):

okay thanks :)

Nnesha (nnesha):

@UnkleRhaukus good work :):)

OpenStudy (unklerhaukus):

\[LHS =(\sin )\left(\tan{\cos }-\cot\cos \right)\\ \qquad=(\sin )\left(\frac{\sin }{\cos }{\cos }-\frac{\cos }{\sin }\cos \right)\\ \qquad=\sin^2-\cos^2\] Now using the the identity \[\sin^2+\cos^2=1\\\qquad\quad\sin^2=1-\cos^2\] \[LHS = (1-\cos^2)-\cos^2\\\qquad=1-2\cos^2\\\qquad=RHS\] (Even without the x's) this looks more messy to me.

OpenStudy (sleepyjess):

Thank you for the help @UnkleRhaukus

ganeshie8 (ganeshie8):

I think so, both are messy in some sense and they are okay for trig proofs. Having said that, the video i attached earlier is talking about a proof like this i guess : \[\begin{align} &\sin^2x + \cos^2x = 1 \\~\\ &\sin^2x -\cos^2 x = 1-2\cos^2x\\~\\ & \sin x \left[\tan x \cos x - \cot x \cos x \right] = 1-2\cos^2x\\~\\ \end{align}\] Here we have started with a known identity and arrived at the desired result. Ofcourse we had to do some bottom up scratch work

OpenStudy (nincompoop):

i like it